# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ Abstract
Pulsars are rapidly rotating neutron stars that emit radiation across the electromagnetic spectrum, from radio to $\gamma$-rays. We use the rapid binary population synthesis suite COMPAS to model the Galactic population of canonical pulsars. We account for both radio and $\gamma$-ray selection effects, as well as the motion of pulsars in the Galactic potential due to natal kicks. We compare our models to the catalogs of pulsars detected in the radio, and those detected in $\gamma$-rays by Fermi, and find broad agreement with both populations. We reproduce the observed ratio of radio-loud to radio-quiet $\gamma$-ray pulsars. We further examine the possibility of low spin-down luminosity ($\dot{E}$ ) pulsars emitting weak, unpulsed $\gamma$-ray emission and attempt to match this with recent stacking results. We demonstrate that the apparent correlation between the latitude of a pulsar and its $\dot{E}$ arises due to natal kicks imparted to pulsars at birth, assuming that all pulsars are born in the Galactic disk.
@@ 1 Introduction
Pulsars are rotating neutron stars most commonly observed by their periodic radio emission. Around 3500 pulsars are known to date (Manchester et al., 2005). Modern radio facilities such as MeerKAT, ASKAP and FAST are uncovering new pulsars (e.g., Han et al., 2021; Padmanabh et al., 2023), whilst modern search techniques are still discovering pulsars in decades-old archival surveys (e.g., Sengar et al., 2023). The ever-growing pulsar population provides an invaluable insight into the final remnants of massive stellar evolution, whilst relativistic binary pulsars allow for exquisitely precise tests of theories of gravity (e.g., Kramer et al., 2021).
@@
Whilst most pulsars are discovered and studied through radio surveys, some are also observed at other wavelengths, including optical (Cocke et al., 1969; Nather et al., 1969), X-rays (Bradt et al., 1969; Fritz et al., 1969; Oegelman et al., 1993) and $\gamma$-rays (Kniffen et al., 1974; Tumer et al., 1984; Thompson et al., 1992). A growing number of pulsars are being detected in $\gamma$-rays by the Large Area Telescope (LAT) onboard the Fermi space telescope (Smith et al., 2023), including some pulsars which were first detected in $\gamma$-rays and then later detected as radio pulsars, a situation exemplified by the Geminga pulsar (Halpern & Holt, 1992; Bertsch et al., 1992; Malofeev & Malov, 1997).
Typically, young pulsars with high spindown luminosity ($\dot{E}$) are found at low Galactic latitudes, and millisecond pulsars (MSPs) are observed in the $\gamma$-rays (e.g., Abdo et al., 2009, 2013; Pletsch et al., 2013; Smith et al., 2023), where $\dot{E}$ is the spin down luminosity given by
$\dot{E}=4\pi^{2}I\dot{P}P^{-3}.$ (1)
@@
where $P$ is the spin period of the pulsar, $\dot{P}$ is its spin derivative and $I$ is its moment of inertia. Comparison of the observed $\gamma$-ray luminosity $L_{\gamma}$ to the spindown luminosity $\dot{E}$ indicates that $\gamma$-ray emission accounts for a large fraction of the energy radiated away by many pulsars (with this fraction $\eta\gtrsim 10\%$; Abdo et al., 2013). Around half of the pulsars detected in $\gamma$-rays are also observed in radio, whilst the other half (typically referred to as radio quiet pulsars) are not (Abdo et al., 2008, 2009, 2013; Smith et al., 2023). The distinction between these two populations likely depends on whether or not their radio emission is beamed towards the Earth (Abdo et al., 2009).
@@
The current sample of $\gamma$-ray detected pulsars is around 300 (Abdo et al., 2013; Smith et al., 2019, 2023) [^1] A list of Fermi detected pulsars is available at https://confluence.slac.stanford.edu/display/GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars and is limited by the sensitivity of Fermi-LAT. There is an apparent ‘death-line’ around a spindown luminosity of $\dot{E}\sim 10^{33}$ erg s-1, which may be a result of the limited sensitivity (Abdo et al., 2013). An extensive search of $\gamma$-ray pulsations from a thousand pulsars (Smith et al., 2019) provides a similar result that the lack of detection of pulsars with lower $\dot{E}$ might be due to the limitation of the current instruments, and suggests population synthesis work to explore this specific subpopulation of pulsars.
@@
Recently, Song et al. (2023) performed a stacking analysis of 12 years of Fermi data, targeting the positions of 362 known radio pulsars with no previous $\gamma$-ray detections. In order to reduce the impact of the Galactic $\gamma$-ray background, they selected pulsars that lie at high galactic latitudes ($|b|>20^{\circ}$). This probed a population of relatively old, low $\dot{E}$ pulsars, different to the usual population of $\gamma$-ray loud pulsars. Song et al. (2023) report the significant detection of $\gamma$-rays from this set of pulsars compared to the expected background. Song et al. (2023) argue that this suggests that the usual $\gamma$-ray death line is predominantly a result of the sensitivity of Fermi. In this paper we attempt to explore the implications of this detection for massive binary and pulsar evolution.
@@
Theoretical population synthesis studies can be used to predict the population of observable pulsars under a set of assumptions. There have been several previous population synthesis studies of the $\gamma$-ray pulsar population. In a seminal work, Faucher-Giguère & Kaspi (2006) performed a population synthesis of the Galactic radio pulsar population. It accounted for almost all of the relevant effects, including radio selection effects, magnetic field decay and pulsar orbits in the Galaxy. Assumptions were made on distributions of pulsar birth spin period and birth magnetic field strength. Since then, there have been a number of similar studies building on this work (e.g., Takata et al., 2011; Dirson et al., 2022) that include radio and $\gamma$-ray selection effect and pulsar evolution theories. Typically, existing pulsar population synthesis works, such as those done by using PsrPopPy (Bates et al., 2014), do not model the details of the full evolution of the binary stellar systems. The study of Kiel et al. (2008) is closest to ours in terms of methodology, but they focus on radio pulsars and do not explore the $\gamma$-ray pulsar population. In this work, we attempt to explore the canonical radio and $\gamma$-ray pulsar populations produced through massive binary stellar evolution. In a follow-up work, we will attempt to understand the full effects on pulsar populations from the stellar evolution physics.
@@
Neutron stars—and hence pulsars—originate from massive stars, and the majority of massive stars are born in binaries (e.g., Sana et al., 2012). Hence, whilst the majority of “normal”, young pulsars (with $P\gtrsim 0.1$ s) are isolated, most of these objects are expected to have been born in binaries, which are subsequently disrupted by the supernova that formed the neutron star. Hence, the binary fraction and spatial distributions (including scale height) of pulsars encode information about their birth locations and natal kicks. It is then important to account for a full picture of stellar and binary evolution for a better understanding of the Galactic pulsar population. Binary population synthesis code suites, such as COMPAS (Riley et al., 2022), are ideally suited for this task, as they are able to rapidly simulate the evolution of large populations of massive binary stars, incorporating all relevant processes such as natal kicks and pulsar evolution. Understanding the properties of the Galactic pulsar population across the radio and $\gamma$-ray wavelengths will aid in constraining pulsar emission physics and the formation and evolution of massive binary stars.
@@
As a first study of the Galactic $\gamma$-ray pulsar population using COMPAS, one of the main purposes of this work is to verify the binary and pulsar physics included in COMPAS. This work and its follow-ups can also be used to make predictions for the population of $\gamma$-ray pulsars that will be observed in the future with continued operation of Fermi-LAT. Other interesting binary pulsar systems include neutron star X-ray binaries (Wijnands & van der Klis, 1998), spider pulsars (Fruchter et al., 1988) and double neutron star binaries (e.g., Hulse & Taylor, 1975; Abbott et al., 2017), and these events can only be studied through the inclusion of binary physics. In follow-up studies, we will investigate known peculiar binary systems that contain at least one neutron star using COMPAS and other suitable tools.
@@ 2.1 COMPAS
We simulate a population of massive binary stars using the rapid binary population synthesis suite COMPAS (Stevenson et al., 2017; Riley et al., 2022). COMPAS is based on the Single Star Evolution (SSE, Hurley et al., 2000) code and the Binary Star Evolution (BSE, Hurley et al., 2002) code with updates and modifications made. Binary interactions such as mass transfer and common envelope evolution are included, but we neglect the effects of tides (Riley et al., 2022).
We use default binary property parameters for our COMPAS simulations (Riley et al., 2022). These parameters and their values/ranges are based on past theoretical and observational studies. The mass of the initially more massive star in the binary is drawn from an initial mass function (IMF, Kroupa, 2001) within the range of 5 to 150 M as we need stars massive enough to leave behind neutron stars at the end of their evolution. The mass ratio ($q=m_{2}/m_{1}$, $m_{2}<m_{1}$) of the binary is drawn from a uniform distribution between 0.01 and 1, with a minimum secondary mass of $0.1$ M (Stevenson et al., 2019). We assume initially circular orbits (Hurley et al., 2002). We draw the orbital separation from a uniform distribution in log space with a minimum of 0.01 and maximum of 1000 AU (Sana et al., 2012). Since we are interested in pulsars formed recently in the Milky Way, we assume a representative value of Galactic metallicity $Z_{\mathrm{Gal}}=0.014$ (Asplund et al., 2009). A list of default COMPAS option values, including mass transfer rate and common envelope evolution, can be found in the link in footnote [^2] https://github.com/FloorBroekgaarden/Double-Compact-Object-Mergers/blob/main/otherFiles/Table_with_detailed_binary__population_synthesis_simulation_settings.pdf.
In this work, neutron stars formed through accretion-induced supernovae, and neutron stars undergoing any recycling/rejuvenation processes are omitted. Only neutron stars formed through core collapse supernovae and electron-capture supernovae that are not in wide, non-interacting binaries (Willcox et al., 2021) are included in the following analysis. The reasoning for this decision is discussed in § 2.2.
@@ 2.2 Neutron star kicks
Neutron stars receive natal kicks at birth (Lyne & Lorimer, 1994; Hobbs et al., 2005; Verbunt et al., 2017; Igoshev, 2020; Igoshev et al., 2021). We use the physical remnant prescription from Mandel & Müller (2020) to relate the remnant mass and kick of a neutron star to the properties of its progenitor. This prescription is based on semi-analytic supernova models, and is calibrated to the Galactic pulsar population (Kapil et al., 2023); specifically, we use $v_{\mathrm{NS}}=520$ km/s for the scaling pre-factor for neutron star kicks, and $\sigma_{\mathrm{NS}}=0.3$ for the fractional stochastic scatter in the kick velocities. Kapil et al. (2023) find the distribution of kicks in single and binary neutron stars to be similar, so we opt to use a single distribution for both. As discussed by Mandel & Igoshev (2023), neutron star kicks inferred from young pulsars may be overestimated by $\sim 15$ if pulsar beams are narrow and preferentially aligned with the kick direction. The Mandel & Müller (2020) prescription is applied equally to both core-collapse supernovae and electron-capture supernovae. We do not allow for electron-capture supernovae to occur in wide, non-interacting binaries, to avoid overproducing low-speed pulsars which are not observed (Willcox et al., 2021).

### v2

@@ Abstract
Pulsars are rapidly rotating neutron stars that emit radiation across the electromagnetic spectrum, from radio to $\gamma$-rays. We use the rapid binary population synthesis suite COMPAS to model the Galactic population of canonical pulsars. We account for both radio and $\gamma$-ray selection effects, as well as the motion of pulsars in the Galactic potential due to natal kicks. We compare our models to the catalogues of pulsars detected in the radio, and those detected in $\gamma$-rays by Fermi, and find broad agreement with both populations. We reproduce the observed ratio of radio-loud to radio-quiet $\gamma$-ray pulsars. We further examine the possibility of low spin-down luminosity ($\dot{E}$ ) pulsars emitting weak, unpulsed $\gamma$-ray emission and attempt to match this with results from a recent $\gamma$-ray stacking survey of these pulsars. We confirm the correlation between the latitude of a pulsar and its $\dot{E}$ arises due to natal kicks imparted to pulsars at birth, assuming that all pulsars are born in the Galactic disk.
@@ 1 Introduction
Pulsars are rotating neutron stars most commonly observed by their periodic radio emission. Around 3500 pulsars are known to date Manchester:2004bp. Modern radio facilities such as MeerKAT, ASKAP and FAST are uncovering new pulsars (Han:2021RAA; Padmanabh:2023arXiv, e.g.,), whilst modern search techniques are still discovering pulsars in decades-old archival surveys (Sengar:2023MNRAS; Sengar:2025MNRAS, e.g.,). The ever-growing pulsar population provides an invaluable insight into the final remnants of massive stellar evolution, whilst relativistic binary pulsars allow for exquisitely precise tests of theories of gravity (Kramer:2021jcw, e.g.,).
@@
Whilst most pulsars are discovered and studied through radio surveys, some are also observed at other wavelengths, including optical Cocke:1969Natur; Nather:1969Natur, X-rays Bradt:1969Natur; Fritz:1969Sci; Oegelman:1993Natur and $\gamma$-rays Kniffen:1974Natur; Tumer:1984Natur; Thompson:1992Natur. A growing number of pulsars are being detected in $\gamma$-rays by the Large Area Telescope (LAT) onboard the Fermi space telescope Fermi-LAT:2023zzt, including some pulsars which were first detected in $\gamma$-rays and then later detected as radio pulsars, a situation exemplified by the Geminga pulsar Halpern:1992Natur; Bertsch:1992Natur; Malofeev:1997Natur.
The spindown luminosity, $\dot{E}$ , is a useful parameter describing a pulsar, and is defined as
$\dot{E}=4\pi^{2}I\dot{P}P^{-3},$ (1)
@@
where $P$ is the spin period of the pulsar, $\dot{P}$ is its spin derivative, and $I$ is its moment of inertia. Most of the pulsars observed in $\gamma$-rays (Abdo:2009Sci; Fermi-LAT:2013svs; Pletsch:2013iva; Fermi-LAT:2023zzt, e.g.,) have high spin down luminosities ($\dot{E}\gtrsim 10^{33}$ erg s-1). This includes both young pulsars (mostly found at low Galactic latitudes) and millisecond pulsars (MSPs). Comparison of the observed $\gamma$-ray luminosity $L_{\gamma}$ to the spindown luminosity $\dot{E}$ indicates that $\gamma$-ray emission accounts for a large fraction of the energy radiated away by many pulsars (Fermi-LAT:2013svs, with this fraction $\eta\gtrsim 10\%$;). On the other hand, comparing the radio luminosity of pulsars to their spindown luminosities suggests that only a small fraction of the energy is radiated away in radio wavelengths. The radio emission efficiency $\xi$, defined as the fraction of rotational energy transformed into radio emission
$\xi=\frac{L_{\textrm{r}}}{\dot{E}},$ (2)
where $L_{\textrm{r}}(\textrm{erg s}^{-1})\approx 7.4\times 10^{27}L_{\textrm{r}}$ (mJy), will always be smaller than 1% Szary2014. Around half of the pulsars detected in $\gamma$-rays are also observed in radio, whilst the other half (typically referred to as radio quiet pulsars) are not Fermi-LAT:2008wkm; Abdo:2009Sci; Fermi-LAT:2013svs; Fermi-LAT:2023zzt. The distinction between these two populations likely depends on whether their radio emission is beamed towards the Earth Abdo:2009Sci.
@@
The current sample of $\gamma$-ray detected pulsars is around 300 Fermi-LAT:2013svs; Smith:2019ApJ; Fermi-LAT:2023zzt [^1] A list of Fermi detected pulsars is available at https://confluence.slac.stanford.edu/display/GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars and is limited by the sensitivity of Fermi-LAT. There is an apparent ‘death-line’ around a spindown luminosity of $\dot{E}\sim 10^{33}$ erg s-1, which may be a result of the limited sensitivity Fermi-LAT:2013svs. An extensive search of $\gamma$-ray pulsations from more than one thousand pulsars Smith:2019ApJ provides a similar result that the lack of detection of pulsars with lower $\dot{E}$ might be due to the limitation of the current instruments, and suggests population synthesis work to explore this specific subpopulation of pulsars.
@@
Recently, Song2023 performed a stacking analysis of 12 years of Fermi data, targeting the positions of 362 known radio pulsars with no previous $\gamma$-ray detections. In order to reduce the impact of the Galactic $\gamma$-ray background, they selected pulsars that lie at high galactic latitudes ($|b|>20^{\circ}$). This probed a population of relatively old, low $\dot{E}$ pulsars, different to the usual population of $\gamma$-ray loud pulsars. Song2023 report the significant detection of $\gamma$-rays from this set of pulsars compared to the expected background. Song2023 argue that this suggests that the usual $\gamma$-ray death-line is predominantly a result of the sensitivity of Fermi. In this paper we attempt to explore the implications of this detection for massive binary and pulsar evolution.
@@
Theoretical population synthesis studies can be used to predict the population of observable pulsars under a set of assumptions. There have been several previous population synthesis studies of the $\gamma$-ray pulsar population. For example, Gonthier:2002ApJ used Monte Carlo simulations to predict the radio and $\gamma$-ray pulsar populations that would be observed by Fermi in the context of the polar cap model for $\gamma$-ray emission. In a seminal work, Faucher-Giguere:2006ApJ performed a population synthesis of the Galactic radio pulsar population. It accounted for almost all of the relevant effects, including radio selection effects, magnetic field decay and pulsar orbits in the Galaxy. Assumptions were made on distributions of pulsar birth spin period and birth magnetic field strength. Since then, there have been a number of similar studies building on this work (Takata2011; Dirson2022, e.g.,) that include radio and $\gamma$-ray selection effects and pulsar evolution theories. Modern pulsar population synthesis studies (Gullon:2014MNRAS; ShiNg:2024ApJ; Graber:2024ApJ; PardoAraujo:2025A&A, e.g.,) make use of updated assumptions regarding the birth distributions, magnetic field decay, and pulsar emission geometry, and some of them use sophisticated machine learning inference techniques to compare model predictions to observations.
@@
Typically, existing pulsar population synthesis works, such as those done by using PsrPopPy Bates:2014MNRAS, do not model the details of the full evolution of the binary stellar systems. Neutron stars—and hence pulsars—originate from massive stars, and the majority of massive stars are born in binaries (Sana:2012Sci; MoeDiStefano::2017ApJS, e.g.,). Hence, whilst the majority of “normal”, young pulsars (with $P\gtrsim 0.1$ s) are isolated, most of these objects are expected to have been born in binaries, which are subsequently disrupted by the supernova that formed the neutron star (BrandtPodsiadlowksi:1995MNRAS; Kalogera:1996ApJ; Renzo:2019A&A, e.g.,). Hence, the binary fraction and spatial distributions (including scale height) of pulsars encode information about their birth locations and natal kicks. It is then important to account for a full picture of stellar and binary evolution for a better understanding of the Galactic pulsar population. Binary population synthesis code suites, such as COMPAS COMPAS, are ideally suited for this task, as they are able to rapidly simulate the evolution of large populations of massive binary stars, incorporating all relevant processes such as natal kicks and pulsar evolution. Understanding the properties of the Galactic pulsar population across the radio and $\gamma$-ray wavelengths will aid in constraining pulsar emission physics and the formation and evolution of massive binary stars.
Some previous studies have combined pulsar population synthesis with binary stellar evolution. The study of Kiel:2008MNRAS is closest to ours in terms of methodology. They use binary population synthesis to study the Galactic pulsar population, but they focus on radio pulsars and do not explore the $\gamma$-ray pulsar population. More recently, Titus:2020MNRAS used binary population synthesis to study the radio pulsar population in the Small Magellanic Cloud. In this work, we attempt to explore the canonical radio and $\gamma$-ray pulsar populations produced through massive binary stellar evolution and compare them to the pulsar populations observed in the Milky Way. In a follow-up work, we will attempt to understand the full effects on pulsar populations from stellar evolution physics.
@@
As a first study of the Galactic $\gamma$-ray pulsar population using COMPAS, one of the main purposes of this work is to verify the binary and pulsar physics included in COMPAS. This work and its follow-ups can also be used to make predictions for the population of $\gamma$-ray pulsars that will be observed in the future with continued operation of Fermi-LAT. Other interesting binary pulsar systems include neutron star X-ray binaries 1998Natur.394..344W, spider pulsars 1988Natur.333..237F and double neutron star binaries (Hulse:1975ApJL; PhysRevLett.119.161101, e.g.,), and these events can only be studied through the inclusion of binary physics. In follow-up studies, we will investigate known peculiar binary systems that contain at least one neutron star using COMPAS and other suitable tools.
@@ 2.1 COMPAS
We simulate a population of massive binary stars using the rapid binary population synthesis suite COMPAS Stevenson:2017tfq; gomez2018; COMPAS; Mandel:2025COMPAS. COMPAS is based on the Single Star Evolution (SSE, SSE,) code and the Binary Star Evolution (BSE, BSE,) code with updates and modifications made as described in COMPAS. Binary interactions such as mass transfer and common envelope evolution are included, but we neglect the effects of tides COMPAS.
We use default binary property parameters for our COMPAS simulations COMPAS. These parameters and their values/ranges are based on past theoretical and observational studies. The mass of the initially more massive star in the binary is drawn from an initial mass function (Kroupa:2000iv, IMF;) within the range of 5 to 150 M as we need stars massive enough to leave behind neutron stars at the end of their evolution. The mass ratio ($q=m_{2}/m_{1}$, $m_{2}<m_{1}$) of the binary is drawn from a uniform distribution between 0.01 and 1, with a minimum secondary mass of $0.1$ M Stevenson:2019ApJ. We assume initially circular orbits BSE. We draw the orbital separation from a uniform distribution in log space with a minimum of 0.01 and maximum of 1000 AU Sana:2012Sci. Since we are interested in pulsars formed recently in the Milky Way, we assume a representative value of Galactic metallicity $Z_{\mathrm{Gal}}=0.014$ Asplund:2009ARA&A. A list of default COMPAS option values, including mass transfer rate and common envelope evolution, can be found in the link in footnote [^2] https://github.com/FloorBroekgaarden/Double-Compact-Object-Mergers/blob/main/otherFiles/Table_with_detailed_binary__population_synthesis_simulation_settings.pdf.
We assume a binary fraction of 100% for massive stars and neglect single stars. This is appropriate for massive O stars ($M>16$ M). For example, Sana:2012Sci find that 70% of massive stars are in close binaries with $\log(P_{\mathrm{orb}}/\mathrm{d})<3.5$, and the binary fraction reaches 100% when including wider binaries Sana:2014ApJS which we include in our COMPAS models (deMinkBelczynski::2015ApJ, see also discussion in). Many of these stars are effectively single, in the sense that they do not interact with their wide binary companion. Assuming a binary fraction of 100% is likely an overestimate for B-type stars, which are the progenitors of most neutron stars and have more typical binary fractions of 70% MoeDiStefano::2017ApJS, see also discussion in Zapartas:2017A&A. However, this effect is likely subdominant to many of the other approximations in this work. We leave a thorough exploration of the binary fraction as a function of mass to future work.
In this work, neutron stars formed through accretion-induced supernovae, and neutron stars undergoing any recycling/rejuvenation processes are omitted. Only neutron stars formed through electron-capture supernovae that are not in wide, non-interacting binaries Willcox:2021kbg, or through core collapse supernovae are included in the following analysis. The reasoning for this decision is discussed in § 2.2.
@@ 2.2 Neutron star kicks
Neutron stars receive natal kicks at birth Lyne:1994Nature; Hobbs:2005MNRAS; Verbunt:2017A&A; Igoshev:2020MNRAS; Igoshev:2021MNRAS; DisbergMandel:2025. It is common in the pulsar population synthesis literature to use a distribution inferred from the velocities of isolated pulsars as the kick distribution (Faucher-Giguere:2006ApJ; Graber:2024ApJ, e.g.,). A common choice is the distribution from Hobbs:2005MNRAS333though see DisbergMandel:2025 who argue that the Maxwellian distribution from Hobbs:2005MNRAS overestimates the typical neutron star velocity due to an error.. However, as we are simulating binary progenitors of pulsars, we need to simulate the underlying kick distribution, not the observed kick distribution. Several classes of supernovae are expected to give rise to reduced kicks, including electron-capture supernovae GessnerJanka:2018ApJ, ultra-stripped supernovae (Tauris:2015MNRAS; Suwa:2015MNRAS; Muller:2019MNRAS, e.g.,) and low-mass iron core-collapse supernovae Stockinger:2020MNRAS; Muller:2019MNRAS.
We use the physical remnant prescription from Mandel:2020qwb to relate the remnant mass and kick of a neutron star to the properties of its progenitor. This prescription is based on semi-analytic supernova models, and is calibrated to the Galactic pulsar population Kapil:2022blf; that is, the velocity distribution of young isolated pulsars matches those inferred from very long baseline interferometry observations Deller:2019ApJ. Specifically, we use $v_{\mathrm{NS}}=520$ km/s for the scaling pre-factor for neutron star kicks, and $\sigma_{\mathrm{NS}}=0.3$ for the fractional stochastic scatter in the kick velocities. Kapil:2022blf find the distribution of kicks in single and binary neutron stars to be similar, so we opt to use a single distribution for both. As discussed by Mandel:2022sxv, neutron star kicks inferred from young pulsars may be overestimated by $\sim 15$ % if pulsar beams are narrow and preferentially aligned with the kick direction. The Mandel:2020qwb prescription is applied equally to both core-collapse supernovae and electron-capture supernovae. We do not allow for electron-capture supernovae to occur in wide, non-interacting binaries, to avoid overproducing low-speed pulsars which are not observed Willcox:2021kbg. This choice is also motivated theoretically as binary interactions are expected to widen the parameter space for the production of electron-capture supernovae compared to single stars (Podsiadlowski:2004ApJ, e.g.,).